OpenStudy (anonymous):

5|3(x + 2)| - 15 = 0

OpenStudy (anonymous):

-1,-3

OpenStudy (anonymous):

|3(x+2)|=15/5=3 3(x+2)=+3,-3 x+2=+1,-1 x=-1,-3